# PIR Over Wireless Channels: Achieving Privacy With Public Responses

### v1

@@ IV-A1 Coding Scheme
According to [15, Lemma 5], there exists a one-to-one mapping function, $\phi(\cdot)$, between a element $\mathbf{s}\in\mathbb{F}_{p}$, and a coset lattice point $\tilde{\bm{\lambda}}\in\Lambda_{f}/\Lambda_{w}$. For simplicity, we assume the coset representatives set is ${\Lambda_{f}\cap{\cal V}(\Lambda_{w})}$, namely:
@@
Without loss of generality, assume $m=1$, i.e., the servers wish to transmit the first symbol from each answer, that is, $a_{1}^{(1)}$ and $a_{2}^{(1)}$. Note that $\mathbf{A}_{1}+\mathbf{A}_{2}$ is equal to either $W_{i}$, or $-W_{i}$. This depends on the sign of $b_{i}$, which is known to the user. In the same way $a_{1}^{(m)}+a_{2}^{(m)}=\pm W_{i}^{(m)}$. To encode the symbols, each server maps the symbols to the codebook, namely, $\bm{\lambda}_{1}=\bm{\phi}(a_{1}^{(1)})$ and $\bm{\lambda}_{2}=\bm{\phi}(a_{2}^{(1)})$. Then, each server samples randomly and independently a lattice point from the relevant coset according to $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$ and $P$ is the transmission power. Equivalently, the servers transmissions are:
@@
where $\alpha=\frac{P}{P+\sigma^{2}_{y}}$ is the MMSE coefficient [14]. The above reduces to the Modulo-Lattice Additive Noise (MLAN) channel ([11]) as follows,
@@
Then, from the AWGN-goodness of $\Lambda_{f}$ and the secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{w}=\frac{\sqrt{P}\sigma_{w}}{\sqrt{P^{2}+\sigma_{w}^{2}}}$), we have [11, 14]:
@@
Thus, combining (22)-(23) together, we obtain Theorem 1.
@@
Although the servers generate $r_{1}$ and $r_{2}$ independently, the user can decode the message. That is because both belong to $\Lambda_{w}$.
Next, we demonstrate that the secrecy-privacy requirement (4) for the $j$-th server is fulfilled.
@@
This lemma follows immediately from the fact that $Q_{g}(\theta)$ is uniformly distributed over $\{0,1\}$ or $\{0,-1\}$ independently of $\theta$ and $W_{1}^{M}$. Additionally, $W_{1}^{M}$ independent with $\theta$ and $Q_{g}(\theta)$. ∎
@@
$\lim_{n\rightarrow\infty}I(\lambda_{1},\lambda_{2};\mathbf{w})=0$ (25)
@@
$\begin{split}&I(\lambda_{1},\lambda_{2};\mathbf{w})\\ &=I(\lambda_{1};\mathbf{w})+I(\lambda_{2};\mathbf{w}\mid\lambda_{1})\\ &=H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w})+H(\lambda_{2}\mid\lambda_{1})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w})\\ &\overset{(a)}{\leq}H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w},\lambda_{2})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w})\\ &\overset{}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{z},\lambda_{2})\\ &\hskip 56.9055pt+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{z},\lambda_{1})\\ &\overset{(b)}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{z})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{z})\\ &\overset{}{=}I(\lambda_{1};\mathbf{z}_{1})+I(\lambda_{2};\mathbf{z}_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
where (a) follows since conditioning reduces entropy and since $\lambda_{1}$ is independent with $\lambda_{2}$. (b) follows again since $\lambda_{1}$ independent with $\lambda_{2}$. The last step follows from the strong secrecy implied by our scheme.
Now, we ready to prove the Secure - Privacy (21) requirement is fulfilled,
$\begin{split}&I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w})\\ &=I(\theta;Q_{j}(\theta),W_{1}^{M})+I(\theta;\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M})\\ &\overset{(a)}{=}H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{j}(\theta),W_{1}^{M},\theta)\end{split}$
(a) from Lemma 1. Assume, without loss of generality, that $j=2$. Thus,
$\begin{split}&H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M},\theta)\\ &=H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}\mid Q_{2}(\theta),W_{1}^{M},\theta,\lambda_{1},\lambda_{2})\\ &\overset{(b)}{\leq}H(\mathbf{w})-H(\mathbf{w}\mid\lambda_{1},\lambda_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
(b) is due to the following Markov chain $\theta\leftrightarrow\left(Q_{1}(\theta),Q_{2}(\theta),W_{1}^{M}\right)\leftrightarrow\lambda_{1},\lambda_{2}\leftrightarrow\mathbf{w}$. The last step is due to Lemma 2. ∎
@@ IV-B A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N>2$ Servers
where $\mathbf{x}_{k}$ represents the signal transmitted by the $k$-th server, and $\mathbf{n}_{y}$ denotes the noise at the user. Similarly, the received signal at the malicious server’s antenna is described as:
$\mathbf{w}(i)=\sum_{k=1}^{N}\mathbf{x}_{k}(i)+\mathbf{n}_{w},$ (27)
@@
As proposed in [5], the user generates two queries (such as (17)) and divides the $N$ servers randomly into two groups of $\left\lfloor{\frac{N}{2}}\right\rfloor$ (in case $N$ is odd we can neglect one server). Each group subsequently constructs its answer based on the query assigned to it. The main distinction from the two-server scheme is the need to ensure that each group generates the same random lattice point (coset representative). To achieve coordination within each group, the user provides each group with a uniform binary vector. This vector is used to generate a sample from the relevant distribution [9] [^2] Generating a sample from a discrete distribution is a well-studied topic [9]. The expected number of fair bits required by an optimal algorithm to generate a random variable $X$ is less than $H(X)+2$., specifically $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$. Thus, the received input at the user and the malicious server is as follows:
@@
Without the loss of generality, assume that the malicious server belongs to group $2$ and can perfectly subtract the answers corresponding to its group (require knowledge about the number of servers participating in the transmission). Its objective becomes decoding $\mathbf{x}_{1}$ from the remaining received signals,
@@
The only thing we have to verify is the decoding stage. Decode: To decode $\mathbf{v}\overset{\Delta}{=}\bm{\phi}(W_{i}^{(1)})$, the user computes the following,
@@ IV-C A Joint Scheme for PIR Over Public Block-Fading AWGN MAC with $N>2$ Servers
In this section, we consider a block-fading AWGN MAC. As discussed in the previous section, the transmissions here are affected not only by noise but also by attenuation due to the channel’s fading characteristics. The core concept of the proposed scheme involves dividing the servers into two groups and exploiting the channel characteristics to cancel the messages between them, leaving only the desired message at the receiver. The primary challenge is ensuring that the responses from both groups are properly aligned. To address this, the group with the higher channel gain reduces its transmission power to match the overall power of the two groups, facilitating alignment at the receiver.
@@ IV-C1 Coding Scheme
$\mathbf{w}(i)=\sum_{k\in\mathcal{S}_{1}}|g_{k}|\mathbf{x}_{1}(i)+\sum_{k\in\mathcal{S}_{2}}|g_{k}|\mathbf{x}_{2}(i)+\mathbf{n}_{w},$ (34)
@@
For the analysis of the curious server, we assume the worst-case scenario: the server has full CSI for the legitimate channels and the inter-server channels ($h_{k}$ and $g_{k}$ for all $k$). Hence, the curious server can cancel part of the messages according to its group. Thus, the curious server observes one of the following, $\mathbf{z}_{1}(i)=\tilde{g}_{1}\mathbf{x}_{1}(i)+\mathbf{n}_{w}$ or $\mathbf{z}_{2}(i)=\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\mathbf{x}_{2}(i)+\mathbf{n}_{w}$. We will take care of both cases. For the first case we choose secrecy-good $\Lambda_{w}^{(n)}$ with respect to
$\tilde{\sigma}_{z_{1}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\frac{1}{\tilde{g}_{1}^{2}}}},$
@@
$\tilde{\sigma}_{z_{2}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\left(\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}\right)^{2}}}.$
Let us define $SNR_{\tilde{y}}=\frac{\tilde{h}_{1}^{2}P}{\sigma_{y}^{2}}$, $SNR_{\tilde{w}_{1}}=\frac{\tilde{g}_{1}^{2}P}{\sigma_{z}^{2}}$, and $SNR_{\tilde{w}_{2}}=\frac{\left(\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\right)^{2}P}{\sigma_{z}^{2}}$.
Thus, from the AWGN-goodness of $\Lambda_{f}$ and the secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{z_{1}}$ and $\tilde{\sigma}_{z_{2}}$), [11, 14], we obtain Theorem 3.
We note that the suggested scheme for the block-fading AWGN channel does not impair the Secure - Privacy (21) requirement, even if the channel coefficients ($\mathbf{h}$ and $\mathbf{g}$) are globally known. Specifically, we have

### v2

@@ IV-A1 Coding Scheme
According to [14, Lemma 5], there exists a one-to-one mapping function, $\phi(\cdot)$, between a element $\mathbf{s}\in\mathbb{F}_{p}$, and a coset lattice point $\tilde{\bm{\lambda}}\in\Lambda_{f}/\Lambda_{w}$. For simplicity, we assume the coset representatives set is ${\Lambda_{f}\cap{\cal V}(\Lambda_{w})}$, namely:
@@
Without loss of generality, assume $m=1$, i.e., the servers wish to transmit the first symbol from each answer, that is, $a_{1}^{(1)}$ and $a_{2}^{(1)}$. Note that $\mathbf{A}_{1}+\mathbf{A}_{2}$ is equal to either $W_{i}$, or $-W_{i}$. This depends on the sign of $b_{i}$, which is known to the user. In the same way $a_{1}^{(m)}+a_{2}^{(m)}=\pm W_{i}^{(m)}$. To encode the symbols, each server maps its symbols to the codebook as follows: $\bm{\lambda}_{1}=\bm{\phi}(a_{1}^{(1)})$ and $\bm{\lambda}_{2}=\bm{\phi}(a_{2}^{(1)})$. Then, each server samples randomly and independently a lattice point from the relevant coset according to $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$ and $P$ is the transmission power. Equivalently, the servers transmissions are:
@@
This construction introduces uncertainty for the curious server, preventing it from inferring the desired message.
@@
where $\alpha=\frac{P}{P+\sigma^{2}_{y}}$ is the MMSE coefficient [13]. Note that the modulo operation is with respect to $\Lambda_{w}$. The above reduces to the Modulo-Lattice Additive Noise (MLAN) channel ([10]) as follows,
@@
Then, from the AWGN-goodness of $\Lambda_{f}$, it follows that $P_{e}$ tends to 00 exponentially fast if:
@@
Additionally, since secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{w}=\frac{\sqrt{P}\sigma_{w}}{\sqrt{P^{2}+\sigma_{w}^{2}}}$), we have [10, 13]:
@@
Combining (23)-(24) together, we obtain Theorem 1.
@@
Although the servers generate $r_{1}$ and $r_{2}$ independently, the user can decode the message.
Next, we demonstrate that the privacy requirement (4) for the $j$-th server is fulfilled. Lemma 1 below shows that regardless of the codewords construction, the query and the messages reveal nothing about the index $\theta$.
@@
This lemma follows since $Q_{g}(\theta)$ is independent of $\theta$. This can be verified from the probability function of $Q(\theta)$, that is $P(Q(\theta)_{j}=a|\theta=i)=P(Q(\theta)_{j}=a)=1/2$ for any $j\in\{1\dots,M)$. Thus, $Q(\theta)$ is uniformly distributed over $\{0,1\}^{M}$ or $\{0,-1\}^{M}$. Additionally, $W_{1}^{M}$ is independent of $\theta$ and $Q_{g}(\theta)$. ∎
@@
$\lim_{n\rightarrow\infty}I(\lambda_{1},\lambda_{2};\mathbf{w}_{j})=0,\quad j\in\{1,2\}$ (26)
@@
$\begin{split}&I(\lambda_{1},\lambda_{2};\mathbf{w}_{j})\\ &=I(\lambda_{1};\mathbf{w}_{j})+I(\lambda_{2};\mathbf{w}_{j}\mid\lambda_{1})\\ &=H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w}_{j})+H(\lambda_{2}\mid\lambda_{1})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w}_{j})\\ &\overset{(a)}{\leq}H(\lambda_{1})-H(\lambda_{1}\mid\mathbf{w}_{j},\lambda_{2})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{1},\mathbf{w}_{j})\\ &\overset{}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{j,w},\lambda_{2})\\ &\hskip 56.9055pt+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{j,w},\lambda_{1})\\ &\overset{(b)}{=}H(\lambda_{1})-H(\lambda_{1}\mid\lambda_{1}+n_{j,w})+H(\lambda_{2})-H(\lambda_{2}\mid\lambda_{2}+n_{j,w})\\ &\overset{}{=}I(\lambda_{1};\mathbf{z}_{1})+I(\lambda_{2};\mathbf{z}_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
where (a) follows since conditioning reduces entropy and since $\lambda_{1}$ is independent with $\lambda_{2}$. (b) follows again since $\lambda_{1}$ independent with $\lambda_{2}$. The final step leverages the strong secrecy ensured by our scheme, specifically preventing the curious server from decoding $\lambda_{i}$ from the received signal.
Now, we ready to prove the Privacy (4) requirement is fulfilled,
$\begin{split}&I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w}_{j})\\ &=I(\theta;Q_{j}(\theta),W_{1}^{M})+I(\theta;\mathbf{w}_{j}\mid Q_{j}(\theta),W_{1}^{M})\\ &\overset{(a)}{=}H(\mathbf{w}_{j}\mid Q_{j}(\theta),W_{1}^{M})-H(\mathbf{w}_{j}\mid Q_{j}(\theta),W_{1}^{M},\theta)\end{split}$
where (a) is by Lemma 1. Assume, without loss of generality, that $j=2$. We have,
$\begin{split}&H(\mathbf{w}_{2}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}_{2}\mid Q_{2}(\theta),W_{1}^{M},\theta)\\ &\leq H(\mathbf{w}_{2}\mid Q_{2}(\theta),W_{1}^{M})-H(\mathbf{w}_{2}\mid Q_{2}(\theta),W_{1}^{M},\theta,\lambda_{1},\lambda_{2})\\ &\overset{(b)}{\leq}H(\mathbf{w}_{2})-H(\mathbf{w}_{2}\mid\lambda_{1},\lambda_{2})\overset{n\rightarrow\infty}{\longrightarrow}0\end{split}$
(b) is due to the following Markov chain $\theta\leftrightarrow\left(Q_{1}(\theta),Q_{2}(\theta),W_{1}^{M}\right)\leftrightarrow\lambda_{1},\lambda_{2}\leftrightarrow\mathbf{w}_{j}$. The last step is due to Lemma 2. ∎
@@ IV-B A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N>2$ Servers
where $\mathbf{x}_{k}$ represents the signal transmitted by the $k$-th server, and $\mathbf{n}_{y}$ denotes the noise at the user. Similarly, the received signal at the curious server antenna is described as:
$\mathbf{w}_{j}(i)=\sum_{k=1}^{N}\mathbf{x}_{k}(i)+\mathbf{n}_{w},$ (28)
@@
As proposed in [5], the user generates two queries (such as (18)) and divides the $N$ servers randomly into two groups of $\left\lfloor{\frac{N}{2}}\right\rfloor$ (in case $N$ is odd we can neglect one server). Each group subsequently constructs its answer based on the query assigned to it. The main distinction from the two-server scheme is lies in ensuring that each group generates the same random lattice point (coset representative). Intuitively, this alignment is crucial to our scheme, as the transmissions must sum coherently to enable successful decoding. To achieve this coordination, the user provides each group with a uniform binary vector, ensuring consistent generation of the lattice points within each group. This vector is used to generate a sample from the relevant distribution [9] [^4] Generating a sample from a discrete distribution is a well-studied topic [9]. The expected number of fair bits required by an optimal algorithm to generate a random variable $X$ is less than $H(X)+2$., specifically $D_{\Lambda_{w}+\bm{\lambda}_{i},\sqrt{P}}$ where $i\in\{1,2\}$. Thus, the received input at the user and the malicious server is as follows:
@@
Without the loss of generality, assume that the malicious server belongs to group $2$ and can perfectly subtract the answers corresponding to its group [^5] Note that this assumption is quite strong, as the malicious server would need to know the exact number of servers participating in the transmission — a nontrivial piece of information.. Its objective becomes decoding $\mathbf{x}_{1}$ from the remaining received signals,
@@
The only thing we have to verify is the decoding stage.
@@ IV-C A Joint Scheme for PIR Over Public Block-Fading AWGN MAC with $N>2$ Servers
In this section, we consider a block-fading AWGN MAC. As discussed in the previous section, the transmissions here are affected not only by noise but also by attenuation due to the channel’s fading characteristics (5). The core concept of the proposed scheme involves dividing the servers into two groups and exploiting the channel characteristics to cancel the messages between them, leaving only the desired message at the receiver. The primary challenge is ensuring that the responses from both groups are properly aligned. To address this, the group with the higher channel gain reduces its transmission power to match the overall power of the two groups, facilitating alignment at the receiver.
@@ IV-C1 Coding Scheme
$\mathbf{w}(i)=\sum_{k\in\mathcal{S}_{1}}g_{k}\mathbf{x}_{1}(i)+\sum_{k\in\mathcal{S}_{2}}g_{k}\mathbf{x}_{2}(i)+\mathbf{n}_{w},$ (35)
@@
Thus, from the AWGN-goodness of $\Lambda_{f}$, $P_{e}$ tends to 00 exponentially fast if
On the other hand, we assume the worst-case scenario for the curious server: that is, the server has full CSI for the legitimate channels and the inter-server channels ($h_{k}$ and $g_{k}$ for all $k$). Additionally, we assume that the curious server knows which servers are participating in the transmission—a nontrivial assumption given that the servers do not communicate with each other. Hence, the curious server can cancel part of the messages according to its group. Thus, the curious server observes one of the following, $\mathbf{z}_{1}(i)=\tilde{g}_{1}\mathbf{x}_{1}(i)+\mathbf{n}_{w}$ or $\mathbf{z}_{2}(i)=\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\mathbf{x}_{2}(i)+\mathbf{n}_{w}$. We will take care of both cases. For the first case we choose $\Lambda_{w}^{(n)}$ with respect to
$\tilde{\sigma}_{w_{1}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\frac{1}{\tilde{g}_{1}^{2}}}},$
@@
$\tilde{\sigma}_{w_{2}}=\frac{\sqrt{P}\sigma_{w}\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}}{\sqrt{P^{2}+\sigma_{w}^{2}\left(\frac{1}{\tilde{g}_{2}}\frac{\tilde{h}_{2}}{\tilde{h}_{1}}\right)^{2}}}.$
Let us define $SNR_{\tilde{y}}=\frac{\tilde{h}_{1}^{2}P}{\sigma_{y}^{2}}$, $SNR_{\tilde{w}_{1}}=\frac{\tilde{g}_{1}^{2}P}{\sigma_{w}^{2}}$, and $SNR_{\tilde{w}_{2}}=\frac{\left(\tilde{g}_{2}\frac{\tilde{h}_{1}}{\tilde{h}_{2}}\right)^{2}P}{\sigma_{w}^{2}}$.
Thus, from secrecy-goodness of $\Lambda_{w}^{(n)}$ (with respect to $\tilde{\sigma}_{w_{1}}$ and $\tilde{\sigma}_{w_{2}}$), [10, 13], we obtain Theorem 3.
We note that the suggested scheme for the block-fading AWGN channel does not impair the Privacy (4) requirement, even if the channel coefficients ($\mathbf{h}$ and $\mathbf{g}$) are globally known. Specifically, we have
@@
To analyze the achievable PIR rate in Theorem 3 we note that the user may choose $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ to maximize:
$\max_{\mathcal{S}_{1},\mathcal{S}_{2}}\left\{\log\left(\min_{i}\left\{\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{w}_{i}}},SNR_{\tilde{y}}\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{y}}}\right\}\right)\right\}$ (39)
Figure 2: The average PIR rate as a function of the number of servers $N$. The transmit power is set to $P=1$ and the different colors represent varying values of $\sigma_{w}$ with $\sigma_{y}$ fixed at $1$.
Finding the optimal solution to this problem is closely related to the subset sum problem, known as NP-hard. Nevertheless, greedy algorithms can be employed to provide suboptimal solutions within a reasonable time frame. In Figure 2 we illustrate the average achievable PIR rate as a function of the number of servers, assuming $\sigma_{y}=\sigma_{w}=1$. Interestingly, a positive PIR rate can still be achieved even with equal noise variances or even greater variance at the user antenna. This is primarily because the channel coefficients between the servers and the user are effectively combined due to the knowledge of CSI. In contrast, the channel coefficients observed by the curious server may cancel each other, leading to a reduced SNR for the curious server.
V Conclusion
We proposed a joint PIR-coding scheme for an AWGN MAC channel, where the public nature of the responses could potentially lead to privacy leakage if one of the servers is malicious. In the block-fading case, we demonstrated that a positive PIR rate is achievable, even when the channel to the malicious server is better than the channel to the user.
It is worth noting that a loss of $1/2$ nat occurs due to using the wiretap coding scheme from [13]. This loss could potentially be avoided by employing the scheme proposed in [15], which utilizes a polar lattice code to eliminate the $\frac{1}{2}$-nat gap from the secrecy capacity of the Gaussian Wiretap Channel (GWC). However, we chose to use the scheme from [13] for the sake of convenience and clarity in presentation, as this choice does not affect the scaling laws of our results.